# Regularity Propagation of Global Weak Solutions to a Navier–Stokes–Cahn–Hilliard System for Incompressible Two-phase Flows with Chemotaxis and Active Transport

### v1

@@ 2.1 Preliminaries
First, we introduce some notations and basic tools that will be used throughout the paper.
[…]
$\|f\|_{L^{q}_{\mathrm{uloc}}(a,+\infty;X)}=\sup_{t\geq a}\|f\|_{L^{q}(t,t+1;X)}<+\infty.$
For $q\in[1,+\infty]$, $W^{1,q}(J;X)$ denotes the space of functions $f$ such that $f\in L^{q}(J;X)$ with $\partial_{t}f\in L^{q}(J;X)$, where $\partial_{t}$ denotes the vector-valued distributional derivative of $f$. For $q=2$, we shall set $H^{1}(J;X)=W^{1,q}(J;X)$ as well as $H^{1}_{\mathrm{uloc}}(a,+\infty;X)=W^{1,q}_{\mathrm{uloc}}(a,+\infty;X)$. $C([a,b];X)$ denotes the Banach space of all continuous functions $f:[a,b]\to X$ equipped with the supremum norm and $BUC([a,+\infty);X)$ denotes the space of all bounded and uniformly continuous functions $f:[a,+\infty)\to X$.
For every $f\in(H^{1}(\Omega))^{\prime}$, we denote its generalized mean over $\Omega$ by $\overline{f}=|\Omega|^{-1}\langle f,1\rangle_{(H^{1})^{\prime},H^{1}}$; if $f\in L^{1}(\Omega)$, then it holds $\overline{f}=|\Omega|^{-1}\int_{\Omega}f\,\mathrm{d}x$. We denote by $L^{2}_{0}(\Omega):=\big{\{}f\in L^{2}(\Omega)\ |\ \overline{f}=0\big{\}}$ the linear subspace of $L^{2}(\Omega)$ with zero mean. Besides, in view of the homogeneous Neumann boundary condition (1.2), we introduce the space $H^{2}_{N}(\Omega):=\big{\{}f\in H^{2}(\Omega)\ |\ \partial_{\bm{n}}f=0\ \textrm{on}\ \partial\Omega\big{\}}$. Then we denote by $\mathcal{A}_{N}\in\mathcal{L}(H^{1}(\Omega),(H^{1}(\Omega))^{\prime})$ the realization of minus Laplacian $-\Delta$ subject to the homogeneous Neumann boundary condition such that
@@
For the linear space $V_{0}=H^{1}(\Omega)\cap L_{0}^{2}(\Omega)$, we denote $V_{0}^{-1}=\big{\{}u\in(H^{1}(\Omega))^{\prime}\ |\ \overline{u}=0\big{\}}\subset(V_{0})^{\prime}$. The restriction of $\mathcal{A}_{N}$ from $V_{0}$ onto $V_{0}^{-1}$ is an isomorphism. In particular, $\mathcal{A}_{N}$ is positively defined on $V_{0}$ and self-adjoint. We denote its inverse map by $\mathcal{N}=\mathcal{A}_{N}^{-1}:V_{0}^{-1}\to V_{0}$. For every $f\in V_{0}^{-1}$, $u=\mathcal{N}f\in V_{0}$ is the unique weak solution of the Neumann problem
@@
For any $f\in V_{0}^{-1}$, we set $\|f\|_{V_{0}^{-1}}=\|\nabla\mathcal{N}f\|$. It is well-known that $f\to\|f\|_{V_{0}^{-1}}$ and $f\to\big{(}\|f-\overline{f}\|_{V_{0}^{-1}}^{2}+|\overline{f}|^{2}\big{)}^{\frac{1}{2}}$ are equivalent norms on $V_{0}^{-1}$ and $(H^{1}(\Omega))^{\prime}$, respectively (see e.g., [41]). From the Poincaré–Wirtinger inequality:
@@
[…]
@@
we see that $\|D\bm{u}\|$ yields an equivalent norm for $\bm{H}^{1}_{0,\mathrm{div}}(\Omega)$. Next, we consider the Stokes operator $\bm{S}=\bm{P}(-\Delta)$ with $D(\bm{S})=\bm{H}^{1}_{0,\mathrm{div}}(\Omega)\cap\bm{H}^{2}(\Omega)$. The space $D(\bm{S})$ is equipped with the inner product $(\bm{S}\bm{u},\bm{S}\bm{v})$ and the equivalent norm $\|\bm{S}\bm{u}\|$ (see e.g., [44, Chapter III]). For any $\bm{u}\in D(\bm{S})$ and $\bm{\zeta}\in\bm{H}^{1}_{0,\mathrm{div}}(\Omega)$, it holds $(\bm{S}\bm{u},\bm{\zeta})=(\nabla\bm{u},\nabla\bm{\zeta})$. The operator $\bm{S}$ is a canonical isomorphism from $\bm{H}^{1}_{0,\mathrm{div}}(\Omega)$ to $(\bm{H}^{1}_{0,\mathrm{div}}(\Omega))^{\prime}$ and we denote its inverse by $\bm{S}^{-1}$. For any $\bm{f}\in(\bm{H}^{1}_{0,\mathrm{div}}(\Omega))^{\prime}$, there is a unique $\bm{u}=\bm{S}^{-1}\bm{f}\in\bm{H}^{1}_{0,\mathrm{div}}(\Omega)$ such that $(\nabla\bm{S}^{-1}\bm{f},\nabla\bm{\zeta})=\langle\bm{f},\bm{\zeta}\rangle_{(\bm{H}^{1}_{0,\mathrm{div}})^{\prime},\bm{H}^{1}_{0,\mathrm{div}}}$ for all $\bm{\zeta}\in\bm{H}^{1}_{0,\mathrm{div}}(\Omega)$. Hence, $\|\nabla\bm{S}^{-1}\bm{f}\|=\langle\bm{f},\bm{S}^{-1}\bm{f}\rangle_{(\bm{H}^{1}_{0,\mathrm{div}})^{\prime},\bm{H}^{1}_{0,\mathrm{div}}}^{\frac{1}{2}}$ yields an equivalent norm on $(\bm{H}^{1}_{0,\mathrm{div}}(\Omega))^{\prime}$.
@@
where the constant $C>0$ only depends on $\Omega$. We also recall the following regularity result for the Stokes operator (see e.g., [44, Chapter III, Theorem 2.1.1] and [23, Lemma B.2]):
@@
Let $\Omega$ be a bounded domain of class $C^{2}$ in $\mathbb{R}^{3}$. For any $\bm{f}\in\bm{L}^{2}_{0,\mathrm{div}}(\Omega)$, there exists a unique pair $(\bm{u},p)\in D(\bm{S})\times V_{0}$ such that $-\Delta\bm{u}+\nabla p=\bm{f}$ a.e. in $\Omega$, that is, $\bm{u}=\bm{S}^{-1}\bm{f}$. Moreover, it holds $\|\bm{u}\|_{\bm{H}^{2}}+\|\nabla p\|\leq C\|\bm{f}\|$, where the constant $C>0$ only depends on $\Omega$.
In the subsequent sections, the symbols $C$, $C_{i}$ stand for generic positive constants that can change from line to line. Specific dependence of these constants in terms of the data will be pointed out if necessary.
@@ 2.2 Main results
    with strictly positive constants $\theta_{0},\ \theta$ satisfying $\theta_{0}-\theta>0$. There exists some small $\epsilon_{0}\in(0,1)$ such that $\varPsi_{0}^{\prime\prime}$ is nondecreasing in $[1-\epsilon_{0},1)$ and nonincreasing in $(-1,-1+\epsilon_{0}]$. Besides, we make the extension $\varPsi_{0}(r)=+\infty$ for $r\notin[-1,1]$.
@@
As indicated in [23, Remark 2.1], one can extend the linear viscosity function (1.4) to $\mathbb{R}$ in such a way to comply (H1). Indeed, since the singular potential guarantees that the phase function satisfies $\varphi\in[-1,1]$, the value of $\nu$ outside of $[-1,1]$ is not important and can be chosen as in (H1). It is straightforward to check that the logarithmic potential (1.5) fulfills the assumption (H2).
We first report a result on the existence of global weak solutions to problem (1.1)–(1.3) on $[0,+\infty)$.
@@
Let $\Omega$ be a bounded domain in $\mathbb{R}^{3}$, with boundary $\partial\Omega$ of class $C^{2}$. Suppose that the hypotheses (H1)–(H3) are satisfied. For any initial data $\bm{v}_{0}\in\bm{L}^{2}_{0,\mathrm{div}}(\Omega)$, $\varphi_{0}\in H^{1}(\Omega)$, $\sigma_{0}\in L^{2}(\Omega)$ with $\|\varphi_{0}\|_{L^{\infty}}\leq 1$ and $|\overline{\varphi}_{0}|<1$, the initial boundary value problem (1.1)–(1.3) admits a global weak solution $(\bm{v},\varphi,\mu,\sigma)$ on $[0,+\infty)$ satisfying
$\bm{v}\in L^{\infty}(0,+\infty;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))\cap L^{2}(0,+\infty;\bm{H}^{1}_{0,\mathrm{div}}(\Omega))\cap W^{1,\frac{4}{3}}_{\mathrm{uloc}}(0,+\infty;(\bm{H}^{1}_{0,\mathrm{div}}(\Omega))^{\prime}),$
$\varphi\in L^{\infty}(0,+\infty;H^{1}(\Omega))\cap L^{4}_{\mathrm{uloc}}(0,+\infty;H^{2}_{N}(\Omega))\cap L^{2}_{\mathrm{uloc}}(0,+\infty;W^{2,6}(\Omega)),$
$\partial_{t}\varphi\in L^{2}_{\mathrm{uloc}}(0,+\infty;(H^{1}(\Omega))^{\prime}),\quad\varPsi^{\prime}(\varphi)\in L_{\mathrm{uloc}}^{2}(0,+\infty;L^{6}(\Omega)),$
$\mu\in L^{2}_{\mathrm{uloc}}(0,+\infty;H^{1}(\Omega)),\quad\nabla\mu\in L^{2}(0,+\infty;\bm{L}^{2}(\Omega)),$
@@
$\sigma\in L^{\infty}(0,+\infty;L^{2}(\Omega))\cap L^{2}_{\mathrm{uloc}}(0,+\infty;H^{1}(\Omega))\cap W^{1,\frac{4}{3}}_{\mathrm{uloc}}(0,+\infty;(H^{1}(\Omega))^{\prime}),$
@@
$\left\langle\partial_{t}\bm{v},\bm{\zeta}\right\rangle_{(\bm{H}^{1}_{0,\mathrm{div}})^{\prime},\bm{H}^{1}_{0,\mathrm{div}}}+\big{(}(\bm{v}\cdot\nabla)\bm{v},\bm{\zeta}\big{)}+\big{(}2\nu(\varphi)D\bm{v},D\bm{\zeta}\big{)}=\big{(}(\mu+\chi\sigma)\nabla\varphi,\bm{\zeta}\big{)},$ (2.1a)
$\left\langle\partial_{t}\varphi,\xi\right\rangle_{(H^{1})^{\prime},H^{1}}+({\bm{v}\cdot\nabla\varphi},\xi)=-(\nabla\mu,\nabla\xi)-\alpha(\overline{\varphi}-c_{0},\xi),$ (2.1b)
$\left\langle\partial_{t}\sigma,\xi\right\rangle_{(H^{1})^{\prime},H^{1}}+({\bm{v}\cdot\nabla\sigma},\xi)+(\nabla\sigma,\nabla\xi)=\chi(\nabla\varphi,\nabla\xi),$ (2.1c)
@@
Let $\Omega$ be a bounded domain in $\mathbb{R}^{3}$, with boundary $\partial\Omega$ of class $C^{3}$. Suppose that the hypotheses (H1)–(H3) are satisfied. Let $(\bm{v},\varphi,\mu,\sigma)$ be a global weak solution to problem (1.1)–(1.3) given by Proposition 2.1. Then the following results hold:
@@
$\varphi\in L^{\infty}(\tau,+\infty;W^{2,6}(\Omega)),\quad\partial_{t}\varphi\in L^{2}(\tau,+\infty;H^{1}(\Omega)),$
$\mu\in L^{\infty}(\tau,+\infty;H^{1}(\Omega))\cap L_{\mathrm{uloc}}^{2}(\tau,+\infty;H^{3}(\Omega)),$
@@
$\sigma\in L^{\infty}(T_{\mathrm{R}},+\infty;H^{1}(\Omega))\cap L^{2}_{\mathrm{uloc}}(T_{\mathrm{R}},+\infty;H^{2}_{N}(\Omega)),\quad\partial_{t}\sigma\in L^{2}_{\mathrm{uloc}}(T_{\mathrm{R}},+\infty;L^{2}(\Omega)).$
@@
Theorem 2.2 (Convergence to Equilibrium).
Let the assumptions of Theorem 2.1 be satisfied. Assume in addition, the potential function $\varPsi$ is real analytic in $(-1,1)$. For every global weak solution $(\bm{v},\varphi,\sigma)$ to problem (1.1)–(1.3), it holds
@@
$-\Delta\varphi_{\infty}+\varPsi^{\prime}(\varphi_{\infty})-\chi\sigma_{\infty}+\beta\mathcal{N}(\varphi_{\infty}-c_{0})=\overline{\varPsi^{\prime}(\varphi_{\infty})}-\chi\overline{\sigma_{\infty}},\quad$ in $\ \text{in }\Omega,$ (2.4a)
$\Delta(\sigma_{\infty}-\chi\varphi_{\infty})=0,$ in $\ \text{in }\Omega,$ (2.4b)
$\partial_{\bm{n}}\varphi_{\infty}=\partial_{\bm{n}}\sigma_{\infty}=0,$ on $\ \text{on }\partial\Omega,$ (2.4c)

### v2

@@ 2.1 Preliminaries
First, we introduce some notation and basic tools that will be used throughout the paper.
Let $X$ be a real Banach space. Its dual space is denoted by $X^{\prime}$, and the duality pairing is denoted by $\langle\cdot,\cdot\rangle_{X^{\prime},X}$. The bold letter $\bm{X}$ denotes the generic space of vectors or matrices, with each component belonging to $X$. We assume that $\Omega\subset\mathbb{R}^{3}$ is a bounded domain with a sufficiently smooth boundary $\partial\Omega$. For any $q\in[1,+\infty]$, $L^{q}(\Omega)$ denotes the Lebesgue space with norm $\|\cdot\|_{L^{q}}$. For $k\in\mathbb{Z}^{+}$, $q\in[1,+\infty]$, $W^{k,q}(\Omega)$ denotes the Sobolev space with norm $\|\cdot\|_{W^{k,q}}$. The Hilbert space $W^{k,2}(\Omega)$ is denoted by $H^{k}(\Omega)$ with the norm $\|\cdot\|_{H^{k}}$. For $s\geq 0$ and $q\in[1,\infty)$, we denote by $H^{s,q}(\Omega)$ the Bessel-potential spaces and by $W^{s,q}(\Omega)$ the Slobodeckij spaces. It holds $H^{s,2}(\Omega)=W^{s,2}(\Omega)$ for all $s$, but for $q\neq 2$ the identity $H^{s,q}(\Omega)=W^{s,q}(\Omega)$ is only true if $s\in\mathbb{N}$. For clarity, we denote by $\mathbb{N}$ the set of natural numbers including zero. If $s\in\mathbb{N}$, then $H^{s,q}(\Omega)$ and $W^{s,q}(\Omega)$ coincide with the usual Sobolev spaces. For simplicity of notation, the norm and inner product on $L^{2}(\Omega)$ (as well as $\bm{L}^{2}(\Omega)$) are denoted by $\|\cdot\|$ and $(\cdot,\cdot)$, respectively. Given a measurable set $J\subset\mathbb{R}$, $L^{q}(J;X)$ with $q\in[1,+\infty]$ denotes the space of Bochner measurable $q$-integrable/essentially bounded functions with values in the Banach space $X$. If $J=(a,b)$ is an interval, we write, for simplicity, $L^{q}(a,b;X)$. Moreover, $f\in L^{q}_{\mathrm{loc}}(a,+\infty;X)$ if and only if $f\in L^{q}(a,b;X)$ for every $b>a$. The space $L^{q}_{\mathrm{uloc}}([a,+\infty);X)$ consists of all measurable functions $f:[a,+\infty)\to X$ such that
$\|f\|_{L^{q}_{\mathrm{uloc}}([a,+\infty);X)}=\sup_{t\geq a}\|f\|_{L^{q}(t,t+1;X)}<+\infty.$
For $q\in[1,+\infty]$, $W^{1,q}(J;X)$ denotes the space of functions $f$ such that $f\in L^{q}(J;X)$ with $\partial_{t}f\in L^{q}(J;X)$, where $\partial_{t}$ denotes the vector-valued distributional derivative of $f$. For $q=2$, we shall set $H^{1}(J;X)=W^{1,q}(J;X)$ as well as $H^{1}_{\mathrm{uloc}}([a,+\infty);X)=W^{1,q}_{\mathrm{uloc}}([a,+\infty);X)$. The notation $C([a,b];X)$ denotes the Banach space of all continuous functions $f:[a,b]\to X$ equipped with the supremum norm. Furthermore, $BC([a,+\infty);X)$ denotes the Banach space of all bounded and continuous functions $f:[a,+\infty)\to X$ endowed with the supremum norm, while $BUC([a,+\infty);X)$ stands for the subspace of all bounded and uniformly continuous functions. In addition, we define $BC_{w}([a,+\infty);X)$ as the topological vector space of all bounded and weakly continuous functions $f:[a,+\infty)\to X$.
For every $f\in(H^{1}(\Omega))^{\prime}$, we denote its generalized mean on $\Omega$ by $\overline{f}=|\Omega|^{-1}\langle f,1\rangle_{(H^{1})^{\prime},H^{1}}$; if $f\in L^{1}(\Omega)$, then it holds $\overline{f}=|\Omega|^{-1}\int_{\Omega}f\,\mathrm{d}x$. We denote by
$L^{2}_{0}(\Omega):=\big\{f\in L^{2}(\Omega)\ |\ \overline{f}=0\big\}$
the linear subspace of $L^{2}(\Omega)$ with zero mean. In addition, in view of the homogeneous Neumann boundary condition (1.2), we introduce the space
$H^{2}_{N}(\Omega):=\big\{f\in H^{2}(\Omega)\ |\ \partial_{\bm{n}}f=0\ \textrm{on}\ \partial\Omega\big\}.$
Then we denote by $\mathcal{A}_{N}\in\mathcal{L}(H^{1}(\Omega),(H^{1}(\Omega))^{\prime})$ the realization of minus Laplacian $\Delta$ subject to the homogeneous Neumann boundary condition such that
@@
For the linear space $V_{0}=H^{1}(\Omega)\cap L_{0}^{2}(\Omega)$, we denote
$V_{0}^{-1}=\big\{u\in(H^{1}(\Omega))^{\prime}\ |\ \overline{u}=0\big\}\subset(V_{0})^{\prime}.$
The restriction of $\mathcal{A}_{N}$ from $V_{0}$ onto $V_{0}^{-1}$ is an isomorphism. In particular, $\mathcal{A}_{N}$ is positively defined on $V_{0}$ and is self-adjoint. We denote its inverse map by $\mathcal{N}=\mathcal{A}_{N}^{-1}:V_{0}^{-1}\to V_{0}$. For every $f\in V_{0}^{-1}$, $u=\mathcal{N}f\in V_{0}$ is the unique weak solution of the Neumann problem
@@
For any $f\in V_{0}^{-1}$, we set $\|f\|_{V_{0}^{-1}}=\|\nabla\mathcal{N}f\|$. It is well-known that $f\to\|f\|_{V_{0}^{-1}}$ and $f\to\big(\|f-\overline{f}\|_{V_{0}^{-1}}^{2}+|\overline{f}|^{2}\big)^{\frac{1}{2}}$ are equivalent norms on $V_{0}^{-1}$ and $(H^{1}(\Omega))^{\prime}$, respectively (see, e.g., [47]). From the Poincaré–Wirtinger inequality:
@@
Next, we introduce the Hilbert spaces of solenoidal vector-valued functions (see, e.g., [19, 51]). We denote by $\bm{L}^{2}_{0,\mathrm{div}}(\Omega)$, $\bm{H}^{1}_{0,\mathrm{div}}(\Omega)$ the closure of
$C_{0,\mathrm{div}}^{\infty}(\Omega;\mathbb{R}^{3})=\big\{\bm{f}\in C_{0}^{\infty}(\Omega;\mathbb{R}^{3}):\,\textrm{div}\bm{f}=0\big\}$
in $\bm{L}^{2}(\Omega)$ and $\bm{H}^{1}(\Omega)$, respectively. For $\bm{L}^{2}_{0,\mathrm{div}}(\Omega)$, we also use $(\cdot,\cdot)$ and $\|\cdot\|$ for its inner product and norm. For any function $\bm{f}\in\bm{L}^{2}(\Omega)$, the Helmholtz–Weyl decomposition holds (see [19, Chapter III]):
@@
we see that $\|D\bm{u}\|$ yields an equivalent norm for $\bm{H}^{1}_{0,\mathrm{div}}(\Omega)$. Next, we consider the Stokes operator $\bm{S}=\bm{P}(-\Delta)$ with $D(\bm{S})=\bm{H}^{1}_{0,\mathrm{div}}(\Omega)\cap\bm{H}^{2}(\Omega)$. The space $D(\bm{S})$ is equipped with the inner product $(\bm{S}\bm{u},\bm{S}\bm{v})$ and the equivalent norm $\|\bm{S}\bm{u}\|$ (see, e.g., [51, Chapter III]). For any $\bm{u}\in D(\bm{S})$ and $\bm{\zeta}\in\bm{H}^{1}_{0,\mathrm{div}}(\Omega)$, it holds $(\bm{S}\bm{u},\bm{\zeta})=(\nabla\bm{u},\nabla\bm{\zeta})$. The operator $\bm{S}$ is a canonical isomorphism from $\bm{H}^{1}_{0,\mathrm{div}}(\Omega)$ to $(\bm{H}^{1}_{0,\mathrm{div}}(\Omega))^{\prime}$ and we denote its inverse by $\bm{S}^{-1}$. For any $\bm{f}\in(\bm{H}^{1}_{0,\mathrm{div}}(\Omega))^{\prime}$, there is a unique $\bm{u}=\bm{S}^{-1}\bm{f}\in\bm{H}^{1}_{0,\mathrm{div}}(\Omega)$ such that
$(\nabla\bm{S}^{-1}\bm{f},\nabla\bm{\zeta})=\langle\bm{f},\bm{\zeta}\rangle_{(\bm{H}^{1}_{0,\mathrm{div}})^{\prime},\bm{H}^{1}_{0,\mathrm{div}}},\quad\forall\,\bm{\zeta}\in\bm{H}^{1}_{0,\mathrm{div}}(\Omega).$
Hence, $\|\nabla\bm{S}^{-1}\bm{f}\|=\langle\bm{f},\bm{S}^{-1}\bm{f}\rangle_{(\bm{H}^{1}_{0,\mathrm{div}})^{\prime},\bm{H}^{1}_{0,\mathrm{div}}}^{\frac{1}{2}}$ yields an equivalent norm on $(\bm{H}^{1}_{0,\mathrm{div}}(\Omega))^{\prime}$.
@@
where the constant $C>0$ depends only on $\Omega$. We also recall the following regularity result for the Stokes operator (see e.g., [51, Chapter III, Theorem 2.1.1] and [28, Lemma B.2]):
@@
Let $\Omega$ be a bounded domain of class $C^{2}$ in $\mathbb{R}^{3}$. For any $\bm{f}\in\bm{L}^{2}_{0,\mathrm{div}}(\Omega)$, there exists a unique pair $(\bm{u},p)\in D(\bm{S})\times V_{0}$ such that $\Delta\bm{u}+\nabla p=\bm{f}$ a.e. in $\Omega$, that is, $\bm{u}=\bm{S}^{-1}\bm{f}$. Moreover, it holds
$\|\bm{u}\|_{\bm{H}^{2}}+\|\nabla p\|\leq C\|\bm{f}\|,$
where the constant $C>0$ only depends on $\Omega$.
In subsequent sections, the symbols $C$, $C_{i}$ represent generic positive constants that can change from line to line. The specific dependence of these constants in terms of the data will be pointed out if necessary.
@@ 2.2 Main results
    with strictly positive constants $\theta_{0},\ \theta$ satisfying $\theta_{0}-\theta>0$. There exists some small $\epsilon_{0}\in(0,1)$ such that $\varPsi_{0}^{\prime\prime}$ is non-decreasing in $1-\epsilon_{0},1)$ and non-increasing in $(-1,-1+\epsilon_{0}]$. Besides, we make the extension $\varPsi_{0}(r)=+\infty$ for $r\notin[-1,1]$.
@@
As indicated in [28, Remark 2.1], one can extend the linear viscosity function (1.4) to $\mathbb{R}$ in such a way to comply (H1). In fact, since the singular potential guarantees that the phase function satisfies $\varphi\in[-1,1]$, the value of $\nu$ outside of [1,1][-1,1] is not important and can be chosen as in (H1). It is straightforward to check that the logarithmic potential (1.5) satisfies the assumption (H2).
We first report a result on the existence of global weak solutions to problem (1.1)–(1.3) on $0,+\infty)$.
@@
Let $\Omega$ be a bounded domain in $\mathbb{R}^{3}$, with boundary $\partial\Omega$ of class $C^{2}$. Suppose that hypotheses (H1)–(H3) are satisfied. For any initial data $\bm{v}_{0}\in\bm{L}^{2}_{0,\mathrm{div}}(\Omega)$, $\varphi_{0}\in H^{1}(\Omega)\cap L^{\infty}(\Omega)$ with $\|\varphi_{0}\|_{L^{\infty}}\leq 1$, $|\overline{\varphi}_{0}|<1$ and $\sigma_{0}\in L^{2}(\Omega)$, the initial boundary value problem (1.1)–(1.3) admits a global weak solution $(\bm{v},\varphi,\mu,\sigma)$ on $0,+\infty)$ satisfying
$\bm{v}\in L^{\infty}(0,+\infty;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))\cap L^{2}(0,+\infty;\bm{H}^{1}_{0,\mathrm{div}}(\Omega))\cap W^{1,\frac{4}{3}}_{\mathrm{uloc}}([0,+\infty);(\bm{H}^{1}_{0,\mathrm{div}}(\Omega))^{\prime}),$
$\varphi\in L^{\infty}(0,+\infty;H^{1}(\Omega))\cap L^{4}_{\mathrm{uloc}}([0,+\infty);H^{2}_{N}(\Omega))\cap L^{2}_{\mathrm{uloc}}([0,+\infty);W^{2,6}(\Omega)),$
$\partial_{t}\varphi\in L^{2}(0,+\infty;(H^{1}(\Omega))^{\prime}),\quad\varPsi^{\prime}(\varphi)\in L_{\mathrm{uloc}}^{2}([0,+\infty);L^{6}(\Omega)),$
$\mu\in L^{2}_{\mathrm{uloc}}([0,+\infty);H^{1}(\Omega)),\quad\nabla\mu\in L^{2}(0,+\infty;\bm{L}^{2}(\Omega)),$
@@
$\sigma\in L^{\infty}(0,+\infty;L^{2}(\Omega))\cap L^{2}_{\mathrm{uloc}}([0,+\infty);H^{1}(\Omega))\cap W^{1,\frac{4}{3}}_{\mathrm{uloc}}([0,+\infty);(H^{1}(\Omega))^{\prime}),$
@@
$\left\langle\partial_{t}\bm{v},\bm{\zeta}\right\rangle_{(\bm{H}^{1}_{0,\mathrm{div}})^{\prime},\bm{H}^{1}_{0,\mathrm{div}}}+\big((\bm{v}\cdot\nabla)\bm{v},\bm{\zeta}\big)+\big(2\nu(\varphi)D\bm{v},D\bm{\zeta}\big)=\big((\mu+\chi\sigma)\nabla\varphi,\bm{\zeta}\big),$ (2.1)
$\left\langle\partial_{t}\varphi,\xi\right\rangle_{(H^{1})^{\prime},H^{1}}+({\bm{v}\cdot\nabla\varphi},\xi)=-(\nabla\mu,\nabla\xi)-\alpha(\overline{\varphi}-c_{0},\xi),$ (2.2)
@@
Remark 2.2.
Applying a uniformly local variant of [6, Theorem 4.10.2, Chapter III] and [2, Lemma 4.1], we find
$\bm{v}\in BC_{w}([0,+\infty);\bm{L}^{2}_{0,\mathrm{div}}(\Omega)),\quad\varphi\in BC_{w}([0,+\infty);H^{1}(\Omega)),\quad\sigma\in BC_{w}([0,+\infty);L^{2}(\Omega)),$
so that the initial data can be attained.
@@
Let $\Omega$ be a bounded domain in $\mathbb{R}^{3}$, with boundary $\partial\Omega$ of class $C^{3}$. Suppose that hypotheses (H1)–(H3) are satisfied. Let $(\bm{v},\varphi,\mu,\sigma)$ be a global weak solution to problem (1.1)–(1.3) given by Proposition 2.1. Then the following results hold:
@@
$\varphi\in L^{\infty}(\tau,+\infty;W^{2,6}(\Omega)),\quad\partial_{t}\varphi\in L^{2}_{\mathrm{uloc}}([\tau,+\infty);H^{1}(\Omega)),$
$\mu\in L^{\infty}(\tau,+\infty;H^{1}(\Omega))\cap L_{\mathrm{uloc}}^{2}([\tau,+\infty);H^{3}(\Omega)),$
@@
$\sigma\in L^{\infty}(T_{\mathrm{R}},+\infty;H^{1}(\Omega))\cap L^{2}_{\mathrm{uloc}}([T_{\mathrm{R}},+\infty);H^{2}_{N}(\Omega)),\quad\partial_{t}\sigma\in L^{2}_{\mathrm{uloc}}([T_{\mathrm{R}},+\infty);L^{2}(\Omega)).$
@@
Theorem 2.2 (Convergence to a single equilibrium).
Let the assumptions of Theorem 2.1 be satisfied. Assume in addition, the potential function $\varPsi$ is real analytic in (1,1)(-1,1). For every global weak solution $(\bm{v},\varphi,\sigma)$ to problem (1.1)–(1.3), it holds
@@
$-\Delta\varphi_{\infty}+\varPsi^{\prime}(\varphi_{\infty})-\chi\sigma_{\infty}+\beta\mathcal{N}(\varphi_{\infty}-c_{0})=\overline{\varPsi^{\prime}(\varphi_{\infty})}-\chi\overline{\sigma_{\infty}},\quad$ $\ \text{in }\Omega,$ (2.6a)
$\Delta(\sigma_{\infty}-\chi\varphi_{\infty})=0,$ $\ \text{in }\Omega,$ (2.6b)
$\partial_{\bm{n}}\varphi_{\infty}=\partial_{\bm{n}}\sigma_{\infty}=0,$ $\ \text{on }\partial\Omega,$ (2.6c)